OpenStudy (anonymous):

8. Over two summers, Ray saved $800.00 and $600.00. The polynomial represents his savings at the beginning of the third year, where x is the growth factor. (The interest rate r is x – 1.) What is the interest rate he needs to save $1,650.00 after the third summer? 800x2 + 600x (1 point) a. 0.1% b. 10.9% c. –285.9% d. 1.1%

OpenStudy (anonymous):

800x^2 + 600x - 1650 = 0, has roots x = 1.1093 r = x - 1 [given from question] r = 1.1093 - 1 = .1093 = 10.93%

OpenStudy (anonymous):

ya? make sense?

OpenStudy (anonymous):

Yea what he said, I overlooked that last part.

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (anonymous):

ur welcome